[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 20, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 21 02:38:27 UTC 2017


HSU Shift Log for Thu Jul 20 2017 22:38:18 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Trying to modify the GTWO for the eastern North pacific on  
Sawfish, the program locked up. A second attempt also failed. Was  
successful instead in modifying GTWO on Tuna. (Except the NAWIPS background  
turned lime green...) (Landsea)

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170721/28d0fc42/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list